OpenStudy (anonymous):

my answer book looks like this: (1−cot2(w)+cos2(w)cot2(w))/csc2(w) down to: (1−cot2(w)(1−cos2(w))/csc2(w) but i'm not sure how

OpenStudy (anonymous):

those were supposed to be squares

OpenStudy (anonymous):

you break it as

OpenStudy (anonymous):

\[\frac{ 1-\cot 2w }{ \csc 2w }+\frac{ \cos 2w \cot2w }{ \csc 2w }\]

OpenStudy (anonymous):

sorry, new to this computerized math part, but they were supposed to be : \[(1-\cot^2(w) + \cos^2(w) * \cot^2(w) ) / \csc^2(w) = \sin^4(w)\]

OpenStudy (anonymous):

but what i cannot understand is how the book demonstrates the following

OpenStudy (anonymous):

\[(1-\cot^2(w) * (1-\cos^2(w)) / \csc^2(w) = \sin^4(w)\]

OpenStudy (anonymous):

i am not sure where the 1-cos^2(w) comes from

OpenStudy (anonymous):

\[\frac{ 1-\cot^2(w) + \cos^2(w) * \cot^2(w) }{ \csc^2(w) } = \sin^4(w)\]

OpenStudy (anonymous):

ok, i misunderstand your cot2 . i got it. i know where it comes from

OpenStudy (anonymous):

sorry i am very new to this today

OpenStudy (anonymous):

wait/ \[\frac{ (1-\cot^2w +\cos^2 w)\cot^2 w }{ \csc^2 w }\]. is it your problem?

OpenStudy (anonymous):

Prove: \[\frac{ 1-\cot^2(w)+\cos^2(w)*\cot^2(w) }{ \csc^2(w) } = \sin^4(w)\]

OpenStudy (anonymous):

ok , i'll try

OpenStudy (anonymous):

ok i rearrange the numerator , just count the numerator only,

OpenStudy (anonymous):

ok i rearrange the numerator , just count the numerator only,\[\cos^2\cot^2 -\cot^2+1 = \cot^2(\cos^2-1)+1= -\cot^2(1-\cos^2)+1\]

OpenStudy (anonymous):

=\[-\cot^2\sin^2 +1 = -\cos^2 +1 =\sin^2\]

OpenStudy (anonymous):

now, add the denominator in, 1/csc^2 = sin^2. you get sin^4. is it right?

OpenStudy (anonymous):

ok, not sure why that was causing me so much difficulty, thank you

OpenStudy (anonymous):

LOl, to smart student, everything is challenged. you think too high. that's it